OpenStudy (anonymous):

please help i have 5 problems that i need to write an equation for in slope intercept form (y=mx+b) for the line that satisfies each set of this condition 1- slope 1/2 passes through (6,4) 2- slope of -2 and passing through the point (-1,-2) 3- passing through the points (6,8) and (9,6) 4- passing through the points (2,3) and (4.9) 5- slope of -2/3 and passing through point (6,-1) If you can help I would loooove it :)

OpenStudy (anonymous):

1. You know the slope and you know two points. Use y=mx+b to find the equation of the line. Plug in x=6 and y=4 and m=1/2 into the equation y=mx+b, then find b. When you have b, you can put it into y=mx+b and there is your equation. 4=(1/2)(6)+b solve for b, then y=(1/2)x+b